OpenStudy (anonymous):

please help Solve the 2nd order homogeneous linear recurrence relation with constant coefficients: f(n+2)-4f(n+1)+4f(n)=0 subject to the start-up conditions f(0)=3; f(1)=4

OpenStudy (anonymous):

hey do you know how to do this?

OpenStudy (anonymous):

if you could be kind enough to show me the work also i would greatly appreciate it...i have no idea how to do this

OpenStudy (anonymous):

I am little rusty, but I think I will figure it out, hold on

OpenStudy (anonymous):

no problem take your time :)

OpenStudy (anonymous):

hi lalaly thaks for coming :)

OpenStudy (anonymous):

thanks*

OpenStudy (lalaly):

no problem:D

OpenStudy (anonymous):

lalaly u left? =/

OpenStudy (anonymous):

: f(n+2)-4f(n+1)+4f(n)=0 use characteristic equation t^2 -4t +4 t=-2,-2 that's repeated root we use the form T(n)= C1 r^n + C2 n r^n C1 -2^n + C2 n -2^n now we have to solve for initial condition C1 -2^0 + C2 *0 -2^0=3 C1=3 C1*-2+C2*-2=4 -6+C2*-2=4 -2 C2=10 C2=-5 T(n)=(3 -2^n + -5( n* -2^n)

OpenStudy (anonymous):

lalaly u left? =/

OpenStudy (anonymous):

so would i have to solve anything after that?

OpenStudy (anonymous):

lalaly u left? =/

OpenStudy (anonymous):

so if i write down every thing you have just written then then the problem should be complete right? the teacher shouldn't ask for more? sorry im just really lost in this class and i gotta pass this take home final to pass the class

OpenStudy (anonymous):

let me check my work

OpenStudy (anonymous):

\[C _{1} r ^{n} + C _{2} n r ^{n}\] is this what u meant to say? ...just trying to figure out where the sub numbers and exponents go

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hey so does it work out? im also asking this question in another post just to see what other people get

OpenStudy (anonymous):

should have been T(n)=(3 2^n + -1( n* 2^n)

OpenStudy (anonymous):

f(n+2)=4f(n+1)-4f(n) f(2)= 4(f1)-4(f0) f0=3 f1=4 f(2)= 16- 12=4 let test it on our equation by plugging in n=2 3(2)^2-1(2)(2)^2 12 -8=4 yes, it checks out